What is the y-intercept of this equation:
y = 2x + 5
DO NOT WRITE AS AN ORDERED PAIR

Found 2 solutions by checkley71, DJ_Sammie:
Answer by checkley71(8403)

USING THE EQUATION FOR A LINE OF Y=mX+b WHERE b=THE Y INTERCEPT THEN b=5 IN THIS EQUATION OF Y=2X+5

Answer by DJ_Sammie(24)

y=mx+b
b=y-int
so the answer is 5